OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that contains the points (4, -7) and (0, 5).

OpenStudy (anonymous):

\[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

Find slope first by using this formula..

OpenStudy (anonymous):

alright is it -31/4 ?

OpenStudy (anonymous):

How will you show your work??

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

how you find the slope just tell me the way..

OpenStudy (anonymous):

I subtraced 5 from -7 and 0 from 4 so I guess

OpenStudy (anonymous):

Hmm I was wrong.. I think, should it be positive?

OpenStudy (anonymous):

What will you get if you subtract 5 from -7..??

OpenStudy (anonymous):

13 because the -- should become a +

OpenStudy (anonymous):

12 or 13??

OpenStudy (anonymous):

Ooops did my math wrong, 12.

OpenStudy (anonymous):

Slope is : -3 Use y - y1 = m(x - x1) m = slope = -3 y1 = 5, x1 = 0 Put to get the equation..